OpenStudy (fionacndg):

Given a geometric sequence in the table below, create the explicit formula and list any restrictions to the domain

OpenStudy (fionacndg):

n an 1 –4 2 20 3 –100 answers: an = –5(–4)^n – 1 where n ≥1 an = –4(–5)^n – 1 where n ≥1 an = –4(5)^n – 1 where n ≥–4 an = 5(–4)^n – 1 where n ≥–4

OpenStudy (fionacndg):

answer is option 2, switch over to peer answer, that is how I got this answer http://www.peeranswer.com/study/mathematics